Papaya Partners is a distributor of papayas. They purchase papayas from individual growers and package them in 10-pound cartons for delivery to their various customers, generally supermarkets. Last month, they budgeted to sell $500,000 worth of cartons at a price of $25 each. Actual sales met a budget of $500,000 at $25 per carton.

Management has received cost information based on actual performance and needs to understand the drivers of the overall variance from the budget. They have asked you, as an analyst in their management accounting department, to calculate and explain the variances. The following data has been provided:

Budget
Cost of fruit @ 10 pounds per carton $ 200,000
Cost of packaging @ 1 pound per carton $ 10,000
Labor costs @ .5 hourse per carton $ 90,000
Total Cost $ 300,000
Actual
Cost of fruit @ 10 pounds per carton $ 244,200
Cost of packaging @ .55 pound per carton $ 11,000
Labor costs @ .75 hourse per carton $ 150,000
Gross Profit $ 405,200
Unfavorable variance    $105,200.00

Specifically, management needs to know the:

  • Standard cost per unit (carton)
  • Actual cost per unit
  • Direct materials price variances
  • Direct materials usage variances
  • Direct labor rate variance
  • Direct labor efficiency variance

In addition, they would like to understand how the variances are calculated and what caused them. They would also like a recommendation on what might be done to improve the variances.

For this assignment, compute all required amounts and explain how the computations were performed. Describe whom you would work with to determine the causes of the variances and hypothesize on what caused the variances. Based on your analysis, recommend actions that management could take to improve the variances.
Answer #1

Explanation:
Let's calculate one-by-one:

1.Standard Cost: It is the budgeted cost of producing one unit of product. The standard cost will be:

Fruit: 200,000/20,000 = 10/carton

Packaging: 10,000/20,000 = 0.5/carton

Labor: 90,000/20,000 = 4.5

Standard Cost = 15/carton (10 + 0.5 + 4.5)

2.Actual cost per unit: It is calculated all the actual costs that are incurred in producing on unit (carton).

Fruit: 244,200/20,000 = 12.21/Carton

Packaging: 11,000/20,000 = 0.55/carton

Labor: 150,000/20,000 = 7.5/carton

Actual cost per carton = 20.26/carton

3.Direct Materials Price Variance: It is the difference between the actual amount spent on direct materials and the standard cost of direct materials.

Standard cost of direct materials: Fruit + Packaging: (200,000 x 1) + (11,000 x 0.50) (standard cost x actual quantity)

Actual amount spent on direct materials Fruit + Packaging : (200,000 x 1.221) + (11,000 x 1) (Actual price x actual quantity)

Variance: 255,200 - 205,500 = 49,700 unfavorable

workings-1: (The information will be used for material price & usage variance)

Fruit:

Standard cost: 20,000 x 10 = 200,000 Pounds; 200,000 Pounds x 1 = $200,000

Standard cost = 1

Standard quantity = 200,000

Actual cost: 20,000 x 10 = 200,000 Pounds; 200,000 x 1.221 = $244,200

Actual cost = 1.221

Actual quantity = 200,000

Packaging:

Standard cost: 20,000 x 1 = 20,000 Pounds x 0.5 = 10,000

Standard rate = 0.5;

Standard quantity = 20,000

Actual cost: 20,000 x 0.55 = 11,000 Pounds x 1 = 11,000

Actual rate = 1

Actual Quantity = 11,000

The variance is unfavorable because the actual spent on purchasing the material is higher than the standard price of direct materials. The purchase department should be consulted in order to determine the reason of unfavorable variance.

4.Direct material usage variance: It is the difference between the actual material usage and budgeted material usage. It is calculated as: (Standard Usage - Actual usage) x Standard cost per unit.

Actual usage: 11,000

Standard Usage: 20,000

Standard price: 0.50

Variance: (20,000 - 11,000) x 0.50 = $4,500 favorable

The usage variance is favorable as actual usage is lower than standard usage, this means that production department is working efficiently and a good quality of packaging in bought by the purchasing department. There is no usage variance in fruit usage as 10 pounds per carton is used under budgeted and actual.

5.Direct Labor rate variance: It is difference between the standard cost of direct labor for standards hours and the standards cost of actual hours. It is calculated as:

Actual Hours x Standard rate : 15,000 x 9 = 135,000

Actual Hours x Actual rate: 15,000 x 10 = 150,000

Variance: 15,000 unfavorable

Standards hours x standard rate:

Production department (usually production supervisor) is to concerned for the inquiries about labor rate variance.

Workings-2: (Will be used for labor rate & efficiency variance)

Standard cost: 20,000 x 0.50 = 10,000 hours; 10,000 x 9 = 90,000

Standard rate = 9/hr

Standard Hours = 10,000

Actual Cost: 20,000 x 0.75 = 15,000 hours; 15,000 x 10 = 150,000

Actual rate = 10/hour

Actual hours = 15,000

6.Direct labor efficiency variance: It is calculated as Actual hours less standard hours multiplied by standard rate.

(Actual Hours - Standard Hours) x standard rate: (15,000 - 10,000) x 9 = 45,000 unfavorable

As ore hours are spent therefore it is an unfavorable variance. The production manager should be contacted in order to determine the reason of the unfavorable variance. Furthermore, the purchase department (usually purchase manager) should also be asked about the quality of material as poor material might take more time.

Conclusion: The labor efficiency variance and direct material variances are the highest variances. The management should look into the production department as both these variances are pointing towards the inefficiency of production department. The management should also inquire about the material quality that is being purchased as poor quality material will affect labor as well as material efficiency. The labor rate unfavorable variance shows another inefficiency of the production department. To conclude, management should take measures to check the quality of the materials purchased and efficiency of the production department; the machinery involved in production should also be checked.